an original investment of $12,000 earns 6.00% interest compounded continuously. what will the investment be worth in 10 years? A=Pe^rt
Answer by stanbon(75887)

an original investment of $12,000 earns 6.00% interest compounded continuously. what will the investment be worth in 10 years?
A(t) = Pe^rt
----
A(10) = 12000*e^(0.06*10) = 12000*e^0.6 = $21865.43
-------------
Cheers,
Stan H.
------------